theyadoreshayyy:

Use the chart below to answer the questions: Country Population (1997 estimate) Area (square km) Indonesia 203,479,000 1,904,569 Denmark 5,248,000 42,077 Ghana 18,339,000 237,533 Which of these countries has the greatest population density? Which has the least? Provide an explanation and proof for your answer to receive full credit.

AZ:

Population density = population / area so you have a table with the populations and the area to calculate the population density, just divide Population density of Indonesia = population / area = 203,479,000 / 1,904,569 = ??? similarly, for Denmark it would be population density = 5,248,000 / 42,077 = ?? and for Ghana population density = 18,339,000 / 237,533 = ??

AZ:

So which one has the highest population density? Which one has the least?